OpenStudy (goformit100):

My( goformit100's ) grandson is about as many days as my son in weeks, and my grandson is as many months as I am in years. My grandson, my son and I together are 120 years. Can you tell me my age in years ?

OpenStudy (goformit100):

@Eyad

OpenStudy (anonymous):

LOL @ "My( goformit100's ) grandson"

OpenStudy (goformit100):

ok ok solve it first

OpenStudy (anonymous):

i hate math o.0 @Eyad this baby seems smart lol

OpenStudy (anonymous):

Let , g = grandson's age s = son's age m = my age ... -My grandson is about as many days as my son is weeks ==> ( s = 7g) . -My grandson is as many months as I am in years.==> m = 12g ==> g + s + m = 120 Substitute for m and s g + 7g + 12g = 120 20g = 120 g =120/20 g = 6 yrs is grandson's age

OpenStudy (anonymous):

@RoxxisaUNICORN : lol